# Unleashing the Potential of Diffusion Models for Incomplete Data Imputation

### v1

@@ 4.2 E-Step: Missing Data Imputation with a Learned Diffusion Model
Specifically, let ${\mathbf{x}}$ be the data to impute, $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$ be the values of observed entries, $\mathbf{m}$ be the $0/1$ indicators of the location of missing entries, and $\tilde{\mathbf{x}}_{t}$ be the imputed data at time $t$. Then, we can obtain the imputed data at time $t-\Delta t$ via combining the observed entries from the forward process on ${\mathbf{x}}_{0}=\mathbf{x}$, and the missing entries from the reverse process on $\tilde{\mathbf{x}}_{t}$ from the prior step $t$ [16, 27]:
@@
Based on the above process, starting from a random noise from the maximum time $T$, i.e., $\tilde{\mathbf{x}}_{T}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$, we are able to obtain a reconstructed $\tilde{\mathbf{x}}_{0}$, such that the observed entries of $\tilde{\mathbf{x}}_{0}$ is the same as those of $\mathbf{x}_{0}$, i.e., $\tilde{\mathbf{x}}^{\rm obs}_{0}=\hat{\mathbf{x}}^{\rm obs}_{0}$. In Theorem 1, we prove that via the algorithm above, the obtained $\tilde{\mathbf{x}}_{0}$ is sampled from the desired conditional distribution, i.e., $\tilde{\mathbf{x}}_{0}\sim p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$.
@@
Let $\tilde{\mathbf{x}}_{T}$ be a sample from the prior distribution $\pi(\mathbf{x})={\mathcal{N}}(\mathbf{0},\sigma^{2}(T)\mathbf{I})$, $\mathbf{x}$ be the data to impute, and the known entries of $\mathbf{x}$ are denoted by $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. The score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is approximated by neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ Applying Eq. 5, Eq. 6, and Eq. 7 iteratively from $t=T$ until $t=0$, then $\tilde{\mathbf{x}}_{0}$ is a sample from $p_{\bm{\theta}}(\mathbf{x})$, under the condition that its observed entries $\tilde{\mathbf{x}}_{0}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. Formally,
@@
See proof in Appendix A.1. Theorem 1 demonstrates that with a learned diffusion model $\bm{\theta}$, we are able to obtain samples exactly from the conditional distribution $p_{\bm{\theta}}(\mathbf{{\bm{x}}}|\mathbf{x}^{\rm obs})$ through the aforementioned imputation process. Therefore, it recovers the E-step in the EM algorithm.
@@
Since the desired imputed data $\mathbf{x}^{{\rm mis}*}$ is the expectation, i.e., $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$. We sample $\tilde{\mathbf{x}}_{0}^{\rm mis}$ for $N$ times, and take the average value as the imputed $\mathbf{x}^{{\rm mis}*}$. The algorithmic illustration of DiffPuter’s E-step is summarized in Algorithm 2.
Input: Data $\mathbf{x}\in p(\mathbf{x})\in{\mathbb{R}}^{1\times d}$.
@@
3# $\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},\mathbf{I})\in{\mathbb{R}}^{d}$
4 while not converging do
@@ 9       Update $\bm{\theta}$ via Adam optimizer
Input: Score network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$. Data with missing values ${\mathbf{x}}\in{\mathbb{R}}^{d}$, mask $\mathbf{m}\in\{0,1\}^{d}$
@@
2
3Sample $\tilde{\mathbf{x}}_{t_{M}}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(t_{M})\mathbf{I})$
@@
6             $\mathbf{x}_{t_{i-1}}^{\rm forward}=\mathbf{x}+\sigma(t_{i-1})\cdot\bm{\varepsilon}$
7             $\mathbf{x}_{t_{i-1}}^{\rm reverse}=\tilde{\mathbf{x}}_{t_{i}}+\int_{t_{i}}^{t_{i-1}}{\rm d}\tilde{\mathbf{x}}_{t_{i}}$
8             $\tilde{\mathbf{x}}_{t_{i-1}}=\mathbf{m}\odot\mathbf{x}_{t_{i-1}}^{\rm forward}+(1-\mathbf{m})\odot\mathbf{x}_{t_{i-1}}^{\rm reverse}$
@@ 6             $\mathbf{x}_{t_{i-1}}^{\rm forward}=\mathbf{x}+\sigma(t_{i-1})\cdot\bm{\varepsilon}$
4.3 Practice of DiffPuter
Input: Data $\mathbf{x}=(\mathbf{x}^{\rm obs},\mathbf{x}^{\rm mis})\sim p(\mathbf{x})$. Number of iterations $K$.
1
Output: Imputed data at $K$-th iteration $\mathbf{x}^{(K)}$
2 # Analog bit encoding
3 $\mathbf{x}\leftarrow{\rm AnaBit}(\mathbf{x})$
4 $\mathbf{x}^{{\rm obs}(0)}\leftarrow{\rm Standardize}(\mathbf{x}^{{\rm obs}})$
5 $\mathbf{x}^{{\rm mis}(0)}=\bm{0}$
6 $\mathbf{x}^{(0)}=(\mathbf{x}^{{\rm obs}(0)},\mathbf{x}^{{\rm mis}(0)})$
7 for $k\leftarrow$ $1$ to $K$ do
8       $\bm{\theta}^{(k)}=$ M-step $(\mathbf{x}^{(k-1)})$
9       $\mathbf{x}^{(k)}=$ E-step $(\mathbf{x}^{(k-1)},\bm{\theta}^{(k)})$
10
11Return $\mathbf{x}^{(K)}=(\mathbf{x}^{{\rm obs}(K)},\mathbf{x}^{{\rm mis}(K)})$
Algorithm 3 DiffPuter for Missing Data Imputation.
Analog bits for Discrete Variables.
A tabular dataset typically encompasses both numerical (i.e., continuous) data and categorical (i.e., discrete) data. However, conventional diffusion models assume continuous Gaussian noise, thereby rendering them inapplicable directly to categorical variables. To this end, for a column of discrete variables, we encode each category into its corresponding binary code based on its index. A discrete variable with $C$ possible values is encoded into a $\lceil\log_{2}C\rceil$-dimensional vector. For example, when $C=14$, and we obtain a discrete value with index $5=2^{2}+2^{0}$, we transform it into a $4$-dimensional vector $[0,1,0,1]$ , which are treated as float variables in computation. When recovering the categorical variable from the imputed encoding, we set the corresponding entries in the vector to $1$ if they exceed $0.5$, otherwise 00. Finally, we convert the binary 0/1 vector to its corresponding categorical value.
Preprocessing and Initialization.
@@ 9       $\mathbf{x}^{(k)}=$ E-step $(\mathbf{x}^{(k-1)},\bm{\theta}^{(k)})$
Since different columns of the tabular data have distinct meanings and might have distinct scales, following existing literature, we compute the mean and standard deviation of each column’s observed values. Then, each column is standardized with zero mean and unit standard deviation. In addition, the execution of the EM algorithm requires the initialized values of missing entries $\mathbf{x}^{\rm mis(0)}$, which might have a huge impact on the model’s convergence. For simplicity, we initialize the missing entries of each column to be the mean of the column’s observed values, equivalent to setting $\mathbf{x}^{\rm mis(0)}=0$ everywhere (since the data has been standardized).
Iterative Training.
@@
To obtain a more accurate estimation of complete data $\mathbf{x}$, DiffPuter iteratively executes the M-step and E-step. To be specific, let $\mathbf{x}^{(k)}=(\mathbf{x}^{{\rm obs}(k)},\mathbf{x}^{{\rm obs}(k)})=(\mathbf{x}^{{\rm obs}},\mathbf{x}^{{\rm obs}(k)})$ be the estimation of complete data at $k$-th iteration, $\bm{\theta}^{(k)}$ be the diffusion model’s parameters at $k$-th iteration, we write $\bm{\theta}^{(k+1)}$ as a function of $\mathbf{x}^{(k)}$, i.e., $\bm{\theta}^{(k+1)}=$ M-step $(\mathbf{x}^{(k)})$, and $\mathbf{x}^{(k+1)}$ as a function of $\bm{\theta}^{(k+1)}$ and $\mathbf{x}^{(k)}$, i.e., $\mathbf{x}^{(k+1)}=$ E-step $(\mathbf{x}^{(k)},\bm{\theta}^{(k+1)})$. Therefore, with the initialized $\mathbf{x}^{(0)}$ and the maximum iteration $K$, we are able to obtain $\mathbf{x}^{(k)}$ from $k=1$ to $K$.
Combining the above designs, we present the overall description of DiffPuter in Algorithm 3.
@@ 5 Experiments
In this section, we conduct experiments to study the efficacy of the proposed DiffPuter in missing data imputation tasks. The code is available at: https://github.com/hengruizhang98/DiffPuter.
@@ 5.1 Experimental Settings
We evaluate the proposed DiffPuter on ten real-world datasets of varying scales that are publicly available. We consider five datasets of only continuous features: California, Letter, Gesture, Magic, and Bean, and five datasets of both continuous and discrete features: Adult, Default, Shoppers, Beijing, and News. The detailed information of these datasets is presented in Appendix C.2. Following previous works [19, 33], we study three missing mechanisms: 1) For missing completely at random (MCAR), the mask for each dimension of every sample is independently sampled from a Bernoulli distribution. 2) For missing at random (MAR), we first select the columns that have no missing values and then employ a logistic model that takes these non-missing columns as input to generate the missing entries in the remaining columns. 3) For MNAR, the masks are generated by a logistic model taking the masked inputs from MCAR. Differences between the three settings are in Appendix C.3. In the main experiments, we set the missing rate as $r=30\%$. For each dataset, we generate $10$ masks according to the missing mechanism and report the mean and standard deviation of the imputing performance. In this section, we only report the performance in the MCAR setting, while the results of the other two settings are in Appendix D.2.
Baselines.
@@
We compare DiffPuter with $16$ powerful imputation methods from different categories: 1) Distribution-matching methods based on optimal transport, including TDM [33] and MOT [19]. 2) Graph-based imputation methods, including GRAPE [31]: a pure bipartite graph-based framework for data imputation, and IGRM [35]: a graph-based imputation method that iteratively reconstructs the friendship network. 3) Iterative methods, including EM with multivariate Gaussian priors [5], MICE [29], MIRACLE [14], SoftImpute [7], and MissForest [28]. 4) Deep generative models, including MIWAE [18], GAIN [30], MCFlow [25], MissDiff [21] and TabCSDI [34]. It is worth noting that MissDiff and TabCSDI are also based on diffusion. We also compare with two recent SOTA imputing methods ReMasker [3] and HyperImpute [8].
Evaluation Protocols.
@@
For each dataset, we use $70\%$ as the training set, and the remaining $30\%$ as the testing set. All methods are trained on the training set. The imputation is applied to both the missing values in the training set and the testing set. Consequently, the imputation of the training set is the ’in-sample’ setting, while imputing the testing set is the ’out-of-sample’ setting. The performance of DiffPuter is evaluated by the divergence between the predicted values and ground-truth values of missing entries. For continuous features, we use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), while for discrete features, we use Accuracy. The implementation details and hyperparameter settings are presented in Appendix C.4.
@@
Figure 2: MCAR: MAE and RMSE in in-sample imputation tasks. A blank column indicates that the method fails or gets out-of-memory for that dataset. DiffPuter outperforms the most competitive baseline method by $8.10\%$ (MAE score) and $5.64\%$ (RMSE score) by average. Comparison with the entire set of baselines is shown in Fig. 7 in Appendix D.1.
5.2 Main Results
Performance in In-Sample Imputation.
We first evaluate DiffPuter’s performance in the in-sample imputation task. Fig. 2 compares the performance of different imputation methods regarding MAE and RMSE. Due to the large number of baselines, we only present $9$ recent methods out of the $16$ baselines in Fig. 2, and we present the full comparison in Fig. 7 in Appendix D.1. We have the following observations: 1) Across all datasets, DiffPuter provides high-quality imputation results, matching the best methods on some datasets and significantly outperforming the second-best methods on the remaining datasets. 2) The other two diffusion-based methods, MissDiff and TabCSDI, fail to give satisfying imputation performance. 3) The basic EM algorithm with Multivariate Gaussian assumptions demonstrates competitive performance on several datasets, indicating that for relatively simple numerical datasets, basic ML algorithms are not inferior to deep learning methods. 4) Compared to predictive imputation methods, DiffPuter exhibits larger standard deviations because predictive imputation is deterministic, whereas DiffPuter requires stochastic sampling.
Fig. 3 further compares the performance regarding Accuracy for discrete columns. Compared to continuous features, DiffPuter performs on par with SOTA methods for categorical features, not demonstrating a significant advantage as it does with continuous features. Overall, for datasets with mixed-type features, DiffPuter performs on par with state-of-the-art methods for their discrete columns while achieving significantly better results for their continuous columns. Therefore, DiffPuter is better at handling the task of missing data imputation for mixed-type datasets.
Figure 3: MCAR: Acc. in in-sample imputation.
Performance in Out-of-Sample Imputation.

### v2

@@ 4.2 E-Step: Missing Data Imputation with a Learned Diffusion Model
Specifically, let ${\mathbf{x}}$ be the data to impute, $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$ be the values of observed entries, $\mathbf{m}$ be the $0/1$ indicators of the location of missing entries, and $\tilde{\mathbf{x}}_{t}$ be the imputed data at time $t$. Then, we can obtain the imputed data at time $t-\Delta t$ via combining the observed entries from the forward process on ${\mathbf{x}}_{0}=\mathbf{x}$, and the missing entries from the reverse process on $\tilde{\mathbf{x}}_{t}$ from the prior step $t$ (Lugmayr et al., 2022; Song et al., 2021b):
@@
Based on the above process, starting at a random noise from the maximum time $T$, i.e., $\tilde{\mathbf{x}}_{T}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(T)\mathbf{I})$, we can obtain a reconstructed $\tilde{\mathbf{x}}_{0}$, such that the observed entries of $\tilde{\mathbf{x}}_{0}$ are the same as those of $\mathbf{x}_{0}$, i.e., $\tilde{\mathbf{x}}^{\rm obs}_{0}=\hat{\mathbf{x}}^{\rm obs}_{0}$. In Theorem 1, we prove that via the algorithm above, the obtained $\tilde{\mathbf{x}}_{0}$ is sampled from the desired conditional distribution, i.e., $\tilde{\mathbf{x}}_{0}\sim p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$.
@@
Let $\tilde{\mathbf{x}}_{T}$ be a sample from the prior distribution $\pi(\mathbf{x})={\mathcal{N}}(\mathbf{0},\sigma^{2}(T)\mathbf{I})$, $\mathbf{x}$ be the data to impute, and the known entries of $\mathbf{x}$ are denoted by $\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. The score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ is approximated by neural network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ Applying Eq. 5, Eq. 6, and Eq. 7 iteratively from $t=T\gg 0$ until $t=0$ with $\Delta t\rightarrow 0$, then $\tilde{\mathbf{x}}_{0}$ is a sample from $p_{\bm{\theta}}(\mathbf{x})$, under the condition that its observed entries $\tilde{\mathbf{x}}_{0}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}$. Formally,
@@
See proof in Appendix B.1. Theorem 1 demonstrates that with a learned diffusion model $\bm{\theta}$, we are able to obtain samples exactly from the conditional distribution $p_{\bm{\theta}}(\mathbf{{\bm{x}}}|\mathbf{x}^{\rm obs})$ through the aforementioned imputation process. For inference, we use Monte Carlo estimation to compute the expectation of the missing values, i.e., $\mathbf{x}^{*}={\mathbb{E}}_{\mathbf{x}}p_{\bm{\theta}}(\mathbf{x}|\mathbf{x}^{\rm obs})\approx\sum_{j=1}^{N}\tilde{\mathbf{x}}_{0}^{(j)}/N$.
@@
Input: Data samples from $p(\mathbf{x})$.
@@
3while not converging do
4       Sample $\mathbf{x}\sim p(\mathbf{x})\in{\mathbb{R}}^{d}$
5       Sample $t\sim p(t)$
6       Sample $\bm{\varepsilon}\sim{\mathcal{N}}(\bm{0},\mathbf{I})\in{\mathbb{R}}^{d}$
@@ 10       Update $\bm{\theta}$ via Adam optimizer
Input: Score network $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$. Data with missing values ${\mathbf{x}}\in{\mathbb{R}}^{d}$, mask $\mathbf{m}\in\{0,1\}^{d}$. Number of samples $N$. Number of sampling steps $M$.
@@
3       Sample $\tilde{\mathbf{x}}_{t_{M}}^{(j)}\sim{\mathcal{N}}(\bm{0},\sigma^{2}(t_{M})\mathbf{I})$
@@
5             $\mathbf{x}_{t_{i-1}}^{{\rm forward},(j)}=\mathbf{x}+\sigma(t_{i-1})\cdot\bm{\varepsilon}$
6             $\mathbf{x}_{t_{i-1}}^{{\rm reverse},(j)}=\tilde{\mathbf{x}}_{t_{i}}^{(j)}+\int_{t_{i}}^{t_{i-1}}{\rm d}\tilde{\mathbf{x}}_{t_{i}}^{(j)}$
7             $\tilde{\mathbf{x}}_{t_{i-1}}^{(j)}=\mathbf{m}\odot\mathbf{x}_{t_{i-1}}^{{\rm forward},(j)}+(1-\mathbf{m})\odot\mathbf{x}_{t_{i-1}}^{{\rm reverse},(j)}$
8
@@ 5             $\mathbf{x}_{t_{i-1}}^{{\rm forward},(j)}=\mathbf{x}+\sigma(t_{i-1})\cdot\bm{\varepsilon}$
Since the desired imputed data $\mathbf{x}^{{\rm mis}*}$ is the expectation, i.e., $\mathbf{x}^{{\rm mis}*}={\mathbb{E}}_{\mathbf{x}^{\rm mis}}p(\mathbf{x}^{\rm mis}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, we sample $\tilde{\mathbf{x}}_{0}^{\rm mis}$ for $N$ times, and take the average value as the imputed $\mathbf{x}^{{\rm mis}*}$. The algorithmic illustration of DiffPuter’s E-step is summarized in Algorithm 2.
4.3 Implementations
Data Processing. Diffusion models are inherently suited for continuous data but not for discrete data. Given that real-world tabular data often contains both types of data, we use one-hot encoding to convert each dimension of discrete data into multi-dimensional 0/1 encoding, thereby treating it as continuous data . After that, we perform standardization on each column of data, ensuring that it has a mean of 0 and a variance of 1. The final imputation performance is measured by the difference between the predicted value and the ground-truth missing value after standardization.
@@ 4.3 Implementations
Initialization of Missing Data. The execution of the EM algorithm requires the initialized values of missing entries $\mathbf{x}^{\rm mis(0)}$, which might have a huge impact on the model’s convergence. For simplicity, we initialize the missing entries of each column to be the mean of the column’s observed values, equivalent to setting $\mathbf{x}^{\rm mis(0)}=0$ everywhere (since the data has been standardized).
@@
Training. To obtain a more accurate estimation of complete data $\mathbf{x}$, DiffPuter iteratively executes the M-step and E-step. To be specific, let $\mathbf{x}^{(k)}=(\mathbf{x}^{{\rm obs}(k)},\mathbf{x}^{{\rm obs}(k)})=(\mathbf{x}^{{\rm obs}},\mathbf{x}^{{\rm obs}(k)})$ be the estimation of complete data at $k$-th iteration, $\bm{\theta}^{(k)}$ be the diffusion model’s parameters at $k$-th iteration, we write $\bm{\theta}^{(k+1)}$ as a function of $\mathbf{x}^{(k)}$, i.e., $\bm{\theta}^{(k+1)}=$ M-step $(\mathbf{x}^{(k)})$, and $\mathbf{x}^{(k+1)}$ as a function of $\bm{\theta}^{(k+1)}$ and $\mathbf{x}^{(k)}$, i.e., $\mathbf{x}^{(k+1)}=$ E-step $(\mathbf{x}^{(k)},\bm{\theta}^{(k+1)})$. Therefore, with the initialized $\mathbf{x}^{(0)}$ and the maximum iteration $K$, we are able to obtain $\mathbf{x}^{(k)}$ from $k=1$ to $K$.
Inference. For in-sample imputation, the imputed values are obtained iteratively during the training process. For out-of-sample imputation, with the observed incomplete data $\mathbf{x}$, the mask $\mathbf{m}$, and the trained score network $\bm{\epsilon}_{\bm{\theta}}$, we directly apply the E-step once in Algorithm 2 to obtain the imputation values $\mathbf{x}^{*}$.
@@ 5 Experiments
In this section, we conduct experiments to study the efficacy of the proposed DiffPuter in missing data imputation tasks.
@@ 5.1 Experimental Settings
Datasets. We evaluate the proposed DiffPuter on ten public real-world datasets of varying scales. We consider five datasets of only continuous features: California, Letter, Gesture, Magic, and Bean, and four datasets of both continuous and discrete features: Adult, Default, Shoppers, and News. The detailed information of these datasets is presented in Appendix D.2. Following previous works (Muzellec et al., 2020; Zhao et al., 2023), we study three missing mechanisms: MCAR, MAR, and MNAR. Differences between the three settings are in Appendix D.3. In this section, we only report the performance in the MCAR setting, while the results of the other two settings are in Appendix E. In the main experiments, we set the missing rate as $r=30\%$. For each dataset, we generate $10$ masks according to the missing mechanism and report the mean and standard deviation of the imputing performance.
@@
Baselines. We compare DiffPuter with $16$ powerful imputation methods from different categories: 1) Distribution-matching methods based on optimal transport, including TDM (Zhao et al., 2023) and MOT (Muzellec et al., 2020). 2) Graph-based imputation methods, including GRAPE (You et al., 2020): a pure bipartite graph-based framework for data imputation, and IGRM (Zhong et al., 2023): a graph-based imputation method that iteratively reconstructs the friendship network. 3) Iterative methods, including EM with multivariate Gaussian priors (García-Laencina et al., 2010), MICE (Van Buuren & Karin, 2011), MIRACLE (Kyono et al., 2021), SoftImpute (Hastie et al., 2015), and MissForest (Stekhoven & Bühlmann, 2012). 4) Deep generative models, including MIWAE (Mattei & Frellsen, 2019), GAIN (Yoon et al., 2018), MCFlow (Richardson et al., 2020), MissDiff (Ouyang et al., 2023) and TabCSDI (Zheng & Charoenphakdee, 2022). It is worth noting that MissDiff and TabCSDI are also based on diffusion. We also compare with two recent SOTA imputing methods ReMasker (Du et al., 2024) and HyperImpute (Jarrett et al., 2022).
@@
Evaluation Protocols. For each dataset, we use $70\%$ as the training set, and the remaining $30\%$ as the testing set. All methods are trained on the training set. The imputation is applied to both the missing values in the training set and the testing set. Consequently, the imputation of the training set is the ’in-sample’ setting, while imputing the testing set is the ’out-of-sample’ setting. The performance of DiffPuter is evaluated by the divergence between the predicted values and ground-truth values of missing entries. For continuous features, we use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), while for discrete features, we use Accuracy. Note that both the MAE and RMSE are calculated based on the input data after standardization (zero mean and unit variance). The implementation details and hyperparameter settings are presented in Appendix D.4.
5.2 Main Results (Mask Completely at Random)
@@
Figure 2: MCAR, In-sample imputation performance on continuous columns: Comparing DiffPuter with 17 baselines on imputing continuous data on all the nine datasets. A blank column indicates that the method fails or gets out-of-memory for that dataset. DiffPuter outperforms the most competitive baseline method by $6.94\%$ (MAE score) and $4.78\%$ (RMSE score) by average. The circled number after the model name denotes its ranking among all methods.
Table 1: MCAR, In-sample imputation performance on discrete columns. Comparison of the imputation accuracy on discrete columns on five datasets. DiffPuter ranks the first among 19 imputation methods. Mean/Median/MF means using the mean/median/most frequent value as the imputation, which will give the same imputation result in this case.
Method Adult Default Shoppers News Average Rank
Statistical
Mean/Median/MF $55.20$±$\pm 0.02$ $53.37$±$\pm 0.11$ $50.60$±$\pm 0.29$ $18.60$±$\pm 0.14$ $44.44$ 15
Traditional iterative
EM (García-Laencina et al., 2010) $61.27$±$\pm 0.12$ $57.80$±$\pm 0.22$ $50.94$±$\pm 0.27$ $39.37$±$\pm 0.06$ $52.35$ 10
MICE (Van Buuren & Karin, 2011) $50.31$±$\pm 0.21$ $51.88$±$\pm 0.30$ $43.43$±$\pm 0.19$ $30.05$±$\pm 1.17$ $43.92$ 16
MIRACLE (Kyono et al., 2021) $62.28$±$\pm 0.21$ $55.79$±$\pm 1.52$ $45.74$±$\pm 0.69$ $39.33$±$\pm 0.27$ $50.79$ 12
SoftImpute (Hastie et al., 2015) $56.18$±$\pm 0.08$ $56.01$±$\pm 0.07$ $50.94$±$\pm 0.27$ $18.70$±$\pm 0.05$ $45.46$ 14
MissForest (Stekhoven & Bühlmann, 2012) $63.51$±$\pm 0.31$ $56.97$±$\pm 0.23$ $51.10$±$\pm 0.61$ $37.55$±$\pm 0.84$ $52.28$ 11
Dist. match
MOT (Muzellec et al., 2020) $63.85$±$\pm 0.16$ $71.95$±$\pm 0.06$ $56.91$±$\pm 0.11$ $38.62$±$\pm 0.20$ $7.835$
TDM (Zhao et al., 2023) $62.64$±$\pm 1.36$ $63.41$±$\pm 7.16$ $55.00$±$\pm 0.46$ $30.62$±$\pm 1.22$ $52.92$ 8
GNN
GRAPE (You et al., 2020) $69.84$±$\pm 0.25$ $73.85$±$\pm 0.01$ $57.32$±$\pm 0.15$ $40.83$±$\pm 0.44$ $60.46$ 3
IGRM (Zhong et al., 2023) $69.21$±$\pm 0.29$ OOM $57.45$±$\pm 0.16$ OOM - -
Generative models
MIWAE (Mattei & Frellsen, 2019) $51.33$±$\pm 0.08$ $48.11$±$\pm 0.02$ $49.23$±$\pm 0.61$ $17.32$±$\pm 0.01$ $41.50$ 18
GAIN (Yoon et al., 2018) $53.73$±$\pm 0.38$ $54.49$±$\pm 1.80$ $47.25$±$\pm 0.95$ $34.20$±$\pm 0.06$ $47.42$ 13
MCFlow (Richardson et al., 2020) $60.31$±$\pm 0.18$ $68.36$±$\pm 0..9$ $52.73$±$\pm 0.29$ $30.08$±$\pm 3.01$ $52.87$ 9
TabCSDI (Zheng & Charoenphakdee, 2022) $58.72$±$\pm 0.25$ $50.57$±$\pm 0.02$ $43.24$±$\pm 0.05$ $17.49$±$\pm 0.14$ $42.51$ 17
MissDiff (Ouyang et al., 2023) $57.39$±$\pm 7.34$ $66.01$±$\pm 1.88$ $48.10$±$\pm 1.86$ $41.39$±$\pm 0.59$ $53.22$ 7
Other
KNN (Pujianto et al., 2019) $63.28$±$\pm 0.52$ $71.64$±$\pm 0.34$ $53.93$±$\pm 0.28$ $38.59$±$\pm 0.17$ $56.86$ 6
Remasker (Du et al., 2024) $69.18$±$\pm 0.08$ $76.88$±$\pm 0.62$ $57.86$±$\pm 0.09$ $44.33$±$\pm 0.08$ $62.06$ 2
HyperImpute (Jarrett et al., 2022) $64.99$±$\pm 0.06$ $74.39$±$\pm 2.07$ 59.19±$\pm 0.04$ $40.29$±$\pm 0.96$ $59.72$ 4
DiffPuter $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}70.12}}$±$\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}\pm 0.17}}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}77.64}}$±$\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}\pm 0.32}}$ $58.82$±${\pm 0.09}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}44.69}}$±$\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}\pm 0.13}}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}62.82}}$ 1
We first evaluate DiffPuter’s performance in the in-sample imputation task. Figure 2 compares the performance of different imputation methods regarding continuous columns using MAE and RMSE. Table 1 compares the performance of different methods regarding discrete columns using accuracy. Our observations on these experimental results are summarized as follows.
Superior performance of DiffPuter. Across all datasets, DiffPuter provides high-quality imputation results, matching the best methods on some datasets and significantly outperforming the second-best methods on the remaining datasets. Compared with the other two diffusion-based methods, MissDiff and TabCSDI, the substantial performance advantage of DiffPuter convincingly demonstrates the correctness and superiority of combining the EM algorithm with Diffusion.
Traditional machine learning methods are still powerful imputors. A well-recognized claim in tabular data machine learning is that traditional machine learning methods can sometimes be better than deep learning methods (Lalande & Doya, 2022; Suh & Song, 2023; Jolicoeur-Martineau et al., 2024), and we have similar observations in missing data imputation. For example, the simple EM algorithm under the mixture-Gaussian assumption and the KNN algorithm give decent imputation performance and outperform a lot of early deep generative imputation methods.
The preference of different types of algorithms for heterogeneous tabular data. Another interesting finding is that discriminative methods (e.g., Remasker, GRAPE, MOT) and generative methods (e.g., DiffPuter) have different preferences for continuous data and discrete data. Overall, generative models appear to be more effective at imputing continuous columns. As shown in Figure 2, our model significantly outperforms these discriminative models. Conversely, discriminative models are more proficient at imputing discrete data: in Table 1, the best discriminative model, Remasker, can achieve results comparable to our method on almost all datasets.